Calculate the wavelengths of Halpha and  Hbeeta lines in Balmer series.

Grade:12

1 Answers

Vikas TU
7 years ago
For Halpha series, n1 = 2 and n2 =3 respectively.
Therfore,
1/k1 =R1^2*(1/4 – 1/9) => 5R/36
k1 = 36/5R
For Hbeta, n1 = 2 and n2 = 4 respectively.
1/k2 = R*1^2(1/4 – 1/16) => 3R/16
k2 = 16/3R respectively.